# An extrapolated and provably convergent algorithm for nonlinear matrix decomposition with the ReLU function

### v1

@@ Abstract
Nonlinear matrix decomposition (NMD) with the ReLU function, denoted ReLU-NMD, is the following problem: given a sparse, nonnegative matrix $X$ and a factorization rank $r$, identify a rank-$r$ matrix $\Theta$ such that $X\approx\max(0,\Theta)$. This decomposition finds application in data compression, matrix completion with entries missing not at random, and manifold learning. The standard ReLU-NMD model minimizes the least squares error, that is, $\|X-\max(0,\Theta)\|_{F}^{2}$. The corresponding optimization problem is nondifferentiable and highly nonconvex. This motivated Saul to propose an alternative model, Latent-ReLU-NMD, where a latent variable $Z$ is introduced and satisfies $\max(0,Z)=X$ while minimizing $\|Z-\Theta\|_{F}^{2}$ (“A nonlinear matrix decomposition for mining the zeros of sparse data”, SIAM J. Math. Data Sci., 2022). Our first contribution is to show that the two formulations may yield different low-rank solutions $\Theta$; in particular, we show that Latent-ReLU-NMD can be ill-posed when ReLU-NMD is not, meaning that there are instances in which the infimum of Latent-ReLU-NMD is not attained while that of ReLU-NMD is. We also consider another alternative model, called 3B-ReLU-NMD, which parameterizes $\Theta=WH$, where $W$ has $r$ columns and $H$ has $r$ rows, allowing one to get rid of the rank constraint in Latent-ReLU-NMD. Our second contribution is to prove the convergence of a block coordinate descent (BCD) applied to 3B-ReLU-NMD and referred to as BCD-NMD. Our third contribution is a novel extrapolated variant of BCD-NMD, dubbed eBCD-NMD, which we prove is also convergent under mild assumptions. We illustrate the significant acceleration effect of eBCD-NMD compared to BCD-NMD, and also show that eBCD-NMD performs well against the state of the art on synthetic and real-world data sets.
@@
Low-rank matrix approximations, nonlinear matrix decomposition, rectified liner unit, block coordinate descent, extrapolation
@@ 1 Introduction
Nonlinear matrix decomposition (NMD) can be defined as follows: given a data matrix $X\in\mathbb{R}^{n\times m}$, a factorization rank $r$, NMD looks for a rank-$r$ matrix $\Theta\in\mathbb{R}^{m\times n}$ such that $X\approx f(\Theta)$, where $f(\Theta)$ applies the scalar function $f$ element-wise on $\Theta$, that is, $[f(\Theta)]_{i,j}=f(\Theta_{i,j})$ for all $i,j$. The choice of $f$ depends on the type of data at hand. For example, one might choose $f(x)=\operatorname{sign}(x)$ for binary data in $\{-1,1\}$ [11], or if $X$ has a multiplicative structure, one can consider the Hadamard product decomposition with $f(x)=x^{2}$, which is also efficient for the compression of dense images  [9, 18] and for representing probabilistic circuits [27]. In this work, we assume $X$ to be nonnegative and sparse, opting for $f$ to be the ReLU function, $f(\cdot)=\max(0,\cdot)$. This NMD decomposition, with $X\approx\max(0,\Theta)$, is referred to as the ReLU decomposition. Intuitively, the ReLU decomposition aims at finding a matrix $\Theta$ that substitutes the zero entries of the target matrix $X$ with negative values in order to decrease its rank. Minimizing the least squares error leads to an optimization problem referred to as ReLU-NMD:
@@
Note that some degree of sparsity on the matrix $X$ is an essential condition for ReLU-NMD to be meaningful in practice. In fact, if $X$ has mostly positive entries, then $\Theta$ will also have mostly positive entries and the ReLU nonlinear activation will be active only for few entries, leading to a model that is similar to standard low-rank matrix approximations such as the truncated singular value decomposition (TSVD). However, it is interesting to note that ReLU-NMD will always lead to a lower error than the TSVD. In fact, let $X_{r}$ be an optimal rank-$r$ approximation of $X\geq 0$ (e.g., obtained with the TSVD), then $\|X-X_{r}\|_{F}^{2}\geq\|X-\max(X_{r},0)\|_{F}^{2}$ since negative entries in $X_{r}$, if any, are replaced by zeros which better approximate the corresponding nonnegative entries in $X$.
ReLU-NMD was introduced by Saul in [34, 33]. It is non-differentiable and non-convex, and hence difficult to tackle. Indeed, to the best of our knowledge, there exists only one algorithm that addresses (1) directly, by parametrizing $\Theta=WH$ and using a coordinate descent approach, optimizing one entry of $W$ and $H$ at a time; see [4]. All other existing approaches rely on Saul’s alternative formulation, dubbed Latent-ReLU-NMD, and defined as
@@
The objective function in (2) is differentiable, while the subproblems in variables $Z$ and $\Theta$ can be solved up to global optimality; see Section 4. However, for a general matrix $X$, the link between the solutions of (1) and (2) has not been explored in the literature. Our first contribution, in Section 3, is to understand better the connection between the two.
A third formulation was proposed in [35] using the parametrization $\Theta=WH$, where $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$, which is common practice for low-rank matrix approximation, as the rank of $\Theta$ does not exceed $r$ [43, 38]. This formulation, referred to as 3B-ReLU-NMD model, is
(3) $\min_{Z,W,H}\lVert Z-WH\rVert_{F}^{2}\quad\mbox{such that}\quad\max(0,Z)=X,$
where 3B stands for 3 blocks of variables. The objective function in (3) is again jointly nonconvex, but it is convex with respect to each matrix variable, $W$, $H$ and $Z$, and can be solved computationally faster, as the rank constraint is not explicitly involved. Problem (3) can be tackled using the block coordinate descent (BCD) method, optimizing each matrix variable sequentially [35, 41].
@@
Section 2 motivates the study of ReLU-NMD by presenting several applications. In Section 3, we clarify the relationship between ReLU-NMD and Latent ReLU-NMD. In particular, we provide a matrix $X$ for which the infimum of the rank-1 Latent-ReLU-NMD is not attained, showing the ill-poseness of Latent-ReLU-NMD (Lemma 3.1). On the other hand, for the same matrix $X$, ReLU-NMD has a non-empty set of optimal rank-1 solutions (Lemma 3.3). This shows that, in general, the two models do not share the same set of optimal low-rank solutions, that is, the sets
@@
do not necessarily coincide. However, we show that any feasible point $(\widehat{Z},\widehat{\Theta})$ of the Latent-ReLU-NMD model satisfies the relation
$\lVert X-\max(0,\widehat{\Theta})\rVert_{F}\leq 2\lVert\widehat{Z}-\widehat{\Theta}\rVert_{F}.$
This guarantees that any approximated solution of Latent-ReLU-NMD provides an approximated solution for ReLU-NMD whose residual is at most twice that of Latent-ReLU-NMD.
In Section 4, we first briefly review the state-of-the-art algorithms for solving Latent-ReLU-NMD and 3B-ReLU-NMD (Section 4.1). We then prove convergence of the BCD scheme for solving 3B-ReLU-NMD (BCD-NMD) (Section 4.2), demonstrating that it falls within the framework of BCD with strictly quasi-convex subproblems [16]. In addition, we introduce a novel extrapolated version of BCD-NMD (eBCD-NMD) (Section 4.3), adapting the well-known LMaFit algorithm for matrix completion [43]. We prove subsequence convergence of eBCD-NMD (Theorem 4.9). To the best of our knowledge, BCD-NMD and eBCD-NMD are the first algorithms for solving 3B-ReLU-NMD with convergence guarantees.
In Section 5, we provide extensive numerical experiments comparing our new approach, eBCD-NMD, with the state of the art. We test the algorithms on various applications: matrix completion with ReLU sampling, Euclidean distance matrix completion (EDMC), compression of sparse data, and low-dimensional embedding. In particular, our experiments show that eBCD-NMD consistently accelerates BCD-NMD and performs favorably with the state of the art, while having stronger theoretical guarantees.
2 Applications of ReLU-NMD
In this section, we first show the link between the ReLU decomposition and neural networks. Then we present some applications of the ReLU decomposition, illustrating how nonlinear models can be used in a variety of different contexts.
2.1 Neural network interpretation
The increasing interest in nonlinear models can be attributed to the rising popularity of neural networks, which use nonlinear functions as activations in their hidden layers. Furthermore, low-rank nonlinear models, involving matrix decompositions, have been studied in the context of neural networks [11, 23, 28, 36]. Let us discuss the link between ReLU-NMD and ReLU networks. Figure 1 illustrates a one hidden-layer neural network.
$g$$f$$W$$Y$$a_{i}$$b_{i}$
Figure 1: One hidden-layer neural network with 5 input and output nodes and 3 hidden nodes.
Let $a_{i}\in\mathbb{R}^{p}$ be $p$-dimensional input and $b_{i}\in\mathbb{R}^{m}$ the corresponding output vectors, for $i=1,\dots,n$. Define the input and output matrices $A=[a_{1},\dots,a_{n}]\in\mathbb{R}^{p\times n}$ and $B=[b_{1},\dots,b_{n}]\in\mathbb{R}^{m\times n}$, respectively. We also denote $Y\in\mathbb{R}^{r\times p}$ the matrix of the weights between the input layer and hidden layer, and $W\in\mathbb{R}^{m\times r}$ the matrix of the weights between the hidden layer and the output layer. The functions $g$ and $f$ are the activation functions for the first and second layer, respectively. We are interested in finding an approximation of a given data $X$ using this neural network with one hidden layer. If we choose the error measure to be the least squares error, we need to minimize
(4) $\lVert X-B\rVert_{F}^{2}\;=\;\lVert X-f(Wg(YA))\rVert_{F}^{2}.$
This shows that the hidden representation of each data point is a nonlinear combination of the inputs of the network, that is, $h_{i}=g(Ya_{i})\in\mathbb{R}^{r}$, or $H=g(YA)\in\mathbb{R}^{r\times n}$ in matrix form. NMD is a relaxation of (4) in which the hidden representation of the input matrix, $H$, can be any real matrix of appropriate dimensions. Furthermore, the number of hidden neurons is equal to the rank of the decomposition, and it needs to be chosen a priori as when designing the structure of the network. If the number of hidden neurons is smaller than the dimension of the inputs, the hidden layer is a low-dimensional embedding of the data.
@@ 2.2 Compression of sparse data
An intuitive application of ReLU decomposition models is the compression of sparse data. Let $X\in\mathbb{R}^{m\times n}$ be a sparse and nonnegative matrix with $\operatorname{nnz}(X)$ non-zero entries, and let $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$ be such that $X\approx\max(0,WH)$. As long as $r(m+n)<\operatorname{nnz}(X)$, storing the factors $W$ and $H$ requires less memory than storing $X$. ReLU models can outperform the standard linear compression techniques such as the TSVD on sparse data, such as sparse images or dictionaries [34, 33, 35, 41, 4]. An interesting example is the identity matrix of any dimension that can be exactly reconstructed using a rank-3 ReLU decomposition [11, 34].
@@ 2.3 Matrix completion with ReLU sampling
The class of matrix completion problems with entries missing not at random (MNR) is characterized by the fact that the probability of an entry being missing depends on the matrix itself. For example, when completing a survey, people are likely to share nonsensitive information such as the name or the surname, while they might be more hesitant to disclose their phone number. Therefore, we might expect the missing entries to be concentrated in some area of the matrix more than others. Even though MNR matrix completion is still a relatively unexplored area, dropping the assumption that missing entries are independent of the matrix is a challenging problem that attracts more and more people as the increasing number of related works testifies [6, 29, 13, 7, 26]. Among them, Liu et al. [26] investigate the matrix completion problem when only positive values are observed, matrix completion with ReLU sampling. This instance of matrix completion falls within the MNR problems since the missing entries depend on the sign of the target matrix. Assume that $X$ is a nonnegative and sparse matrix and that there exists a rank-$r$ matrix $\Theta$ such that $X=\max(0,\Theta)$. In this case, finding a ReLU decomposition is equivalent to recovering the negative values, observing only the positive ones; thus, it is equivalent to matrix completion with ReLU sampling. This implies that we can use the algorithms for computing a ReLU decomposition to address this matrix completion problem, as illustrated in [26].
Inspired by the connection with ReLU sampling matrix completion, we introduce an equivalent formulation of 3B-ReLU-NMD that we will use throughout the paper. Let $\Omega=\{(i,j)\ |\ X_{ij}>0\}$ be the set of positive entries of $X$ and let $P_{\Omega}(X)$ be equal to $X$ in $\Omega$ and zero elsewhere. We rewrite 3B-ReLU-NMD in (3) as
(5) $\operatorname{argmin}_{Z,W,H}\frac{1}{2}\lVert Z-WH\rVert_{F}^{2}\quad\mbox{such that}\quad P_{\Omega}(Z)=P_{\Omega}(X),\quad P_{\Omega^{C}}(Z)\leq 0,$
where $\Omega^{C}$ is the complement of $\Omega$ and a reformulation of the constraint $\max(0,Z)=X$ is introduced. Using the matrix factorization $\Theta=WH$ is a standard approach in matrix completion [5, 19, 20, 21, 38], see [30, 25] for detailed surveys on the topic. To our knowledge, the problem (5), without the constraint $P_{\Omega^{C}}(Z)\leq 0$, was first addressed in [43] in the context of matrix completion. The novelties in our model are the additional information that the missing entries have negative signs, which translates into the constraints $P_{\Omega^{C}}(Z)\leq 0$, and the explicit dependence of the known index set $\Omega$ from the target matrix $X$. Since the formulations (3) and (5) are equivalent, we will refer to both as 3B-ReLU-NMD.
@@ 2.4 Euclidean distance matrix completion (EDMC)
We provide now one practical example of the MNR matrix completion problem that can be solved using the ReLU decomposition. It is the recovery of a low-rank matrix for which only the smallest or the largest entries are observed. Let $\Theta$ be a low-rank matrix, and assume that we observe all the entries in $\Theta$ smaller than a known threshold $d$. We want to recover the missing entries, which are the ones larger than $d$. We can model this problem using a rank-1 modified ReLU decomposition: $X=\max(0,dee^{T}-\Theta)$, where $e$ denotes the vector of all ones of appropriate dimension. The matrix $X$ is nonzero exactly in the entries where $\Theta$ is smaller than $d$. Of course, an analogous model can be derived when only the largest entries are observed. The 3B-ReLU-NMD formulation in (5) can be easily adapted to include the rank-1 modification, that is,
(6) $\operatorname{argmin}_{Z,W,H}\frac{1}{2}\lVert dee^{T}-WH-Z\rVert_{F}^{2}\quad\mbox{such that}\quad P_{\Omega}(Z)=P_{\Omega}(X),\quad P_{\Omega^{C}}(Z)\leq 0.$
The same idea applies also to the Latent-ReLU-NMD and all the algorithms can be adjusted to take into account the additional rank-one term. This formulation is particularly interesting when the matrix $\Theta$ contains the squared distances between a collection of points $p_{1},\dots,p_{n}\in\mathbb{R}^{d}$. This problem is often referred to as Euclidean distance matrix completion (EDMC) [1, 12, 22, 39], and is particularly meaningful in sensor network localization problems where sensors can only communicate with nearby sensors. Let $P=[p_{1},\dots,p_{n}]\in\mathbb{R}^{d\times n}$, then the corresponding Euclidean distance matrix is $\Theta_{i,j}=\|p_{i}-p_{j}\|_{2}^{2}$ for all $i,j$, so that $\Theta=e\text{diag}(P^{T}P)^{T}+\text{diag}(P^{T}P)e^{T}+2P^{T}P$ and it has at most rank $r=d+2$ and thus the rank of the factorization in (6) is known in advance. We will show in Section 5 that this slightly modifed ReLU-NMD model allows us to solve EDMC in this context.

### v2

# package/hyperref/before

@@ Abstract
ReLU matrix decomposition (RMD) is the following problem: given a sparse, nonnegative matrix $X$ and a factorization rank $r$, identify a rank-$r$ matrix $\Theta$ such that $X\approx\max(0,\Theta)$. RMD is a particular instance of nonlinear matrix decomposition (NMD) that finds application in data compression, matrix completion with entries missing not at random, and manifold learning. The standard RMD model minimizes the least squares error, that is, $\|X-\max(0,\Theta)\|_{F}^{2}$. The corresponding optimization problem, least-squares RMD (LS-RMD), is nondifferentiable and highly nonconvex. This motivated Saul to propose an alternative model, dubbed Latent-RMD, where a latent variable $Z$ is introduced and satisfies $\max(0,Z)=X$ while minimizing $\|Z-\Theta\|_{F}^{2}$ (“A nonlinear matrix decomposition for mining the zeros of sparse data”, SIAM J. Math. Data Sci., 2022). Our first contribution is to show that the two formulations may yield different low-rank solutions $\Theta$. We then consider a reparametrization of the Latent-RMD, called 3B-RMD, in which $\Theta$ is substituted by a low-rank product $WH$, where $W$ has $r$ columns and $H$ has $r$ rows. Our second contribution is to prove the convergence of a block coordinate descent (BCD) approach applied to 3B-RMD. Our third contribution is a novel extrapolated variant of BCD, dubbed eBCD, which we prove is also convergent under mild assumptions. We illustrate the significant acceleration effect of eBCD compared to BCD, and also show that eBCD performs well against the state of the art on synthetic and real-world data sets.
@@
Low-rank matrix approximations, nonlinear matrix decomposition, rectified linear unit, block coordinate descent, extrapolation.
@@ 1 Introduction
Nonlinear matrix decomposition (NMD) can be defined as follows: given a data matrix $X\in\mathbb{R}^{n\times m}$, a factorization rank $r$, NMD looks for a rank-$r$ matrix $\Theta\in\mathbb{R}^{m\times n}$ such that $X\approx f(\Theta)$, where $f(\Theta)$ applies the scalar function $f$ element-wise on $\Theta$, that is, $f(\Theta)]_{i,j}=f(\Theta_{i,j})$ for all $i,j$. The choice of $f$ depends on the type of data at hand. For example, one might choose $f(x)=\operatorname{sign}(x)$ for binary data in $\{-1,1\}$ [delsarte1989low], or $f(x)=x^{2}$ for nonnegative data with a multiplicative structure, resulting in the Hadamard product decomposition which is efficient for the compression of dense images [ciaperoni2024hadamard, hyeon2021fedpara, wertz2025efficient] and for the representation of probabilistic circuits [loconte2024subtractive]. For a more detailed description of the choices for the function $f$; see [awari2025alternating]. In this work, we assume $X$ to be nonnegative and sparse, opting for $f$ to be the ReLU function, $f(\cdot)=\max(0,\cdot)$. This NMD, with $X\approx\max(0,\Theta)$, is referred to as the ReLU matrix decomposition (RMD). This model has a close connection with a two-layer neural network, using ReLU as the activation function in the output layer; see [saul2022geometrical, saul2022nonlinear] for more details. Intuitively, RMD aims at finding a matrix $\Theta$ that substitutes the zero entries of the target matrix $X$ with negative values in order to decrease its rank. Note that some degree of sparsity on the matrix $X$ is an essential condition for RMD to be meaningful in practice. In fact, if $X$ has mostly positive entries, then $\Theta$ will also have mostly positive entries, and the ReLU nonlinear activation will be active only for few entries, leading to a model that is similar to the truncated singular value decomposition (TSVD).
Minimizing the least squares error leads to an optimization problem referred to as LS-RMD:
@@
It is interesting to note that LS-RMD will always lead to a lower error than the TSVD. In fact, let $X_{r}$ be an optimal rank-$r$ approximation of $X\geq 0$ (e.g., obtained with the TSVD), then
$\|X-X_{r}\|_{F}^{2}\geq\|X-\max(X_{r},0)\|_{F}^{2}$ since negative entries in $X_{r}$, if any, are replaced by zeros which better approximate the corresponding nonnegative entries in $X$.
LS-RMD was introduced by Saul in [saul2022geometrical, saul2022nonlinear]. It is non-differentiable and non-convex, and hence difficult to tackle. Indeed, to the best of our knowledge, there exist only a few algorithms that address (1) directly [awari2025alternating, awaricoordinate]. All other existing approaches rely on Saul’s alternative formulation, dubbed Latent-RMD, and defined as
@@
The objective function in (2) is differentiable, while the subproblems in the variables $Z$ and $\Theta$ can be solved up to global optimality; see Section 4. However, for a general matrix $X$, the link between the solutions of (1) and (2) has not been explored in the literature. Our first contribution, in Section 3, is to better understand the connection between the two.
If we reparametrize $\Theta=WH$ in (2), where $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$, as suggested in [seraghiti2023accelerated], we obtain the following equivalent problem:
(3) $\min_{Z,W,H}\lVert Z-WH\rVert_{F}^{2}\quad\text{ such that }\quad\max(0,Z)=X.$
We denote this reformulation 3B-RMD, where 3B stands for 3 blocks of variables. The objective function in (3) is again jointly nonconvex, but it is convex with respect to each matrix variable, $W$, $H$ and $Z$, making it more friendly from an algorithmic point of view, as the rank constraint is not explicitly involved. Problems (2) and (3) can be tackled using the block coordinate descent (BCD) method, optimizing each matrix variable alternatively [seraghiti2023accelerated, wang2024momentum].
@@
Section 2 motivates the study of RMD by presenting several applications. In Section 3, we clarify the relationship between LS-RMD and Latent-RMD. In particular, we provide a matrix $X$ for which the infimum of the rank-1 Latent-RMD is not attained, showing the ill-poseness of this model ( Example 3.1). On the other hand, for the same matrix $X$, LS-RMD has a non-empty set of optimal rank-1 solutions ( Example 3.3). This shows that, in general, the two models do not share the same set of optimal low-rank solutions, that is, the sets
@@
do not necessarily coincide. However, we show that any feasible point $(\bar{Z},\bar{\Theta})$ of the Latent-RMD model satisfies the relation
[…]
This guarantees that any approximated solution of Latent-RMD provides an approximated solution for LS-RMD whose residual is at most four times that of Latent-RMD.
In Section 4, we first briefly review the state-of-the-art algorithms for solving Latent-RMD and 3B-RMD (Section 4.1). We then prove the convergence of the BCD scheme for solving 3B-RMD (Section 4.2), demonstrating that it falls within the framework of BCD with strictly quasi-convex subproblems [grippo2000convergence]. In addition, we introduce a novel extrapolated version of BCD ( eBCD) (Section 4.3), adapting the well-known LMaFit algorithm for matrix completion [wen2012solving]. We prove the subsequence convergence of eBCD (Theorem 4.9). To the best of our knowledge, BCD and eBCD are the first algorithms for solving 3B-RMD with convergence guarantees.
In Section 5, we provide extensive numerical experiments comparing our new approach, eBCD, with the state of the art. We test the algorithms on various applications: matrix completion with ReLU sampling, Euclidean distance matrix completion (EDMC), compression of sparse data, and low-dimensional embedding. In particular, our experiments show that eBCD consistently accelerates BCD and performs favorably compared to the state of the art, while having stronger theoretical guarantees.
2 Applications of RMD
In this section, we present some applications of RMD, illustrating how nonlinear models can be used in a variety of contexts.
@@ 2.1 Compression of sparse data
The first application is the compression of sparse data. Let $X\in\mathbb{R}^{m\times n}$ be a sparse and nonnegative matrix with $\operatorname{nnz}(X)$ non-zero entries, and let $W\in\mathbb{R}^{m\times r}$ and $H\in\mathbb{R}^{r\times n}$ be such that $X\approx\max(0,WH)$. As long as $r(m+n)<\operatorname{nnz}(X)$, storing the factors $W$ and $H$ requires less memory than storing $X$. RMD typically outperforms standard linear compression techniques, including the TSVD on sparse data, such as sparse images or dictionaries [awaricoordinate, saul2022geometrical, saul2022nonlinear, seraghiti2023accelerated, wang2024momentum]. An interesting example is the identity matrix of any dimension, which can be exactly reconstructed using a rank-3 RMD [delsarte1989low, saul2022nonlinear].
@@ 2.2 Matrix completion with ReLU sampling
The class of matrix completion problems with entries missing not at random (MNR) is characterized by the probability of an entry being missing that depends on the matrix itself. For example, when completing a survey, people are likely to share nonsensitive information such as their name or surname, while they might be more hesitant to disclose their phone number. Therefore, we might expect the missing entries to be concentrated in some areas of the matrix more than others. Even though MNR matrix completion is still a relatively unexplored area, dropping the assumption that missing entries are independent of the matrix is a challenging problem that attracts more and more researchers as the increasing number of related works testifies [bhattacharya2022matrix, bi2021absence, ganti2015matrix, liu2024symmetric, naik2022truncated]. Among them, Liu et al. [liu2024symmetric] investigate the matrix completion problem when only positive values are observed, namely matrix completion with ReLU sampling. This instance of matrix completion falls within the MNR problems since the missing entries depend on the sign of the target matrix. Assume that $X$ is a nonnegative and sparse matrix and that there exists a rank-$r$ matrix $\Theta$ such that $X=\max(0,\Theta)$. In this case, finding an RMD is equivalent to recovering the negative values observing the positive ones; thus, it is equivalent to matrix completion with ReLU sampling. This implies that we can use the algorithms for RMD to address this matrix completion problem, as illustrated in [liu2024symmetric].
Inspired by the connection to ReLU sampling matrix completion, we introduce an equivalent formulation of 3B-RMD that we will use throughout the paper. Let $\Omega=\{(i,j)\ |\ X_{ij}>0\}$ be the set of positive entries of $X$, and let $P_{\Omega}(X)$ be equal to $X$ in $\Omega$ and zero elsewhere. We rewrite 3B-RMD in (3) as
(4) $\min_{Z,W,H}\frac{1}{2}\lVert Z-WH\rVert_{F}^{2}\quad\mbox{such that}\quad P_{\Omega}(Z)=P_{\Omega}(X),\quad P_{\Omega^{C}}(Z)\leq 0,$
where $\Omega^{C}$ is the complement of $\Omega$ and a reformulation of the constraint $\max(0,Z)=X$ is introduced. To our knowledge, the problem (4), without the constraint $P_{\Omega^{C}}(Z)\leq 0$, was first addressed in [wen2012solving] in the context of matrix completion. The novelties in our model are the additional information that the missing entries have negative signs, which translates into the constraints $P_{\Omega^{C}}(Z)\leq 0$, and the explicit dependence of the known index set $\Omega$ on the target matrix $X$. Since the formulations (3) and (4) are equivalent, we will refer to both as 3B-RMD.
@@ 2.3 Euclidean distance matrix completion (EDMC)
We provide one practical example of MNR matrix completion problem that can be solved using RMD. Specifically, we address the recovery of a low-rank matrix for which only the smallest or the largest entries are observed. Let $\Theta$ be a low-rank matrix, and assume that we observe all the entries in $\Theta$ smaller than a known threshold $d$. We want to recover the missing entries, which are the ones larger than $d$. We can model this problem using a rank-1 modified RMD: $X=\max(0,dee^{T}-\Theta)$, where $e$ denotes the vector of all ones of appropriate dimension. The matrix $X$ is nonzero exactly in the entries where $\Theta$ is smaller than $d$. Of course, an analogous model can be derived when only the largest entries are observed. The 3B-RMD formulation in (4) can be easily adapted to include the rank-1 modification:
(5) $\min_{Z,W,H}\frac{1}{2}\lVert dee^{T}-WH-Z\rVert_{F}^{2}\quad\mbox{such that}\quad P_{\Omega}(Z)=P_{\Omega}(X),\quad P_{\Omega^{C}}(Z)\leq 0.$
The same idea applies also to the Latent-RMD and all the algorithms can be adjusted to take into account the additional rank-one term. This formulation is particularly interesting when the matrix $\Theta$ contains the squared distances between a collection of points $p_{1},\dots,p_{n}\in\mathbb{R}^{d}$. This problem is often referred to as Euclidean distance matrix completion (EDMC) [alfakih1999solving, fang2012Euclidean, krislock2012euclidean, tasissa2018exact], and is particularly meaningful in sensor network localization problems where sensors can only communicate with nearby sensors. Let $P=[p_{1},\dots,p_{n}]\in\mathbb{R}^{d\times n}$, then the corresponding Euclidean distance matrix is $\Theta_{i,j}=\|p_{i}-p_{j}\|_{2}^{2}$ for all $i,j$, so that $\Theta=e\text{diag}(P^{T}P)^{T}+\text{diag}(P^{T}P)e^{T}+2P^{T}P$ and it has at most rank $r=d+2$ and thus the rank of the factorization in (5) is known in advance. We will show in Section 5 that this slightly modified RMD allows us to solve EDMC in this context.